OpenStudy (gavin39):

Solve 4 over x plus 2 equals 2 over x for x and determine if the solution is extraneous or not. x = -2, extraneous x = -2, non-extraneous x = 2, extraneous x = 2, non-extraneous

OpenStudy (solomonzelman):

\[\frac{4}{x+2}=\frac{2}{x}~~~~~~~~~~~THIS~?\]

OpenStudy (gavin39):

Yes

OpenStudy (solomonzelman):

\[\frac{4}{x+2}=\frac{2}{x}\]if this is true, then \[\frac{x+2}{4}=\frac{x}{2}\]must also be true AGREE?

OpenStudy (gavin39):

Yes

OpenStudy (solomonzelman):

both sides times 4.\[x+2=2x\]\[2=2x-x\]\[2=x\]MAKES SENSE?

OpenStudy (gavin39):

Yes it does!

OpenStudy (solomonzelman):

Is it extraneous? lets plug it in. \[\frac{4}{\color{red} {2}+2}=\frac{2}{\color{red} {2}}\]\[\frac{4}{4}=\frac{2}{2}\]So, yes, it is extraneous.

OpenStudy (gavin39):

ok awesome really appreciate it!

OpenStudy (solomonzelman):

Anytime, gavin39 !